OpenStudy (anonymous):

Determine the [ H3O+] and [OH-] of 1.0 M NaCl if the pH is 6.04

OpenStudy (anonymous):

Since your pH=6.04 Therefore, you can easily solve for the [H3O+] or [H+] (since both are exactly the same thing) using the antilog, or raising 10 to the negative power of the pH. \[[H3O+]=10^{-pH}=10^{-6.04}=9.12011\times10^{-7}M\] Next, to solve for the OH- concentration, remember that: \[[OH-]=Kw \div[H3O+]\] -------OR------- \[[OH-]=10^{-(14-pH)}\] \[remember Kw=1\times10^{-14}\] Kw is the ionization constant of water, which basically shows that(ALWAYS) \[[H3O+]\times[OH-]=Kw=1\times10^{-14}\] So, \[[OH-]=1\times10^{-14}\div(9.12011\times10^{-11}M)=1.09648\times10^{-8}M\] \[[OH-]=10^{-(14-6.04)}=1.90648\times10^{-8}M\] I solved for [OH-] two different ways, but still got the same answer, which basically shows I was doing the same equation but with different steps. Hope this helps!